### v1

# Geometry-Aware Approaches for Balancing Performance and Theoretical Guarantees in Linear Bandits

Abstract
This paper is motivated by recent developments in the linear bandit literature, which have revealed a discrepancy between the promising empirical performance of algorithms such as Thompson sampling and Greedy, when compared to their pessimistic theoretical regret bounds. The challenge arises from the fact that while these algorithms may perform poorly in certain problem instances, they generally excel in typical instances. To address this, we propose a new data-driven technique that tracks the geometry of the uncertainty ellipsoid, enabling us to establish an instance-dependent frequentist regret bound for a broad class of algorithms, including Greedy, OFUL, and Thompson sampling. This result empowers us to identify and “course-correct” instances in which the base algorithms perform poorly. The course-corrected algorithms achieve the minimax optimal regret of order $\widetilde{\mathcal{O}}(d\sqrt{T})$, while retaining most of the desirable properties of the base algorithms. We present simulation results to validate our findings and compare the performance of our algorithms with the baselines.
1 Introduction
Multi-armed bandits (MABs) have garnered significant attention as they provide a well-defined framework and techniques for investigating the exploration-exploitation trade-off in sequential decision-making problems. In the context of MAB problems, a decision-maker sequentially selects actions from a given set and observes corresponding rewards. This paper specifically focuses on the stochastic linear bandit (LB) model, which encompasses MAB and contextual bandit problems as special cases. In this model, the system parameter $\theta^{\star}$ represents an unknown vector in $\mathbb{R}^{d}$, while the actions, also vectors in $\mathbb{R}^{d}$, yield rewards equal to the inner product of $\theta^{\star}$ and the chosen action, along with mean-zero, subgaussian noise. The objective of a policy is to maximize the cumulative reward based on the observed data up to the decision time. The policy’s performance is measured by the cumulative regret, which quantifies the difference between the total rewards achieved by the policy and the maximum achievable reward. Achieving this objective necessitates striking a balance between exploration and exploitation. In the context of LB, this entails selecting actions that aid in estimating the true parameter $\theta^{\star}$ accurately while obtaining optimal rewards. Various algorithms based on the Optimism in the Face of Uncertainty (OFU) principle have been developed to address this challenge, wherein the optimal action is chosen based on the upper confidence bound (UCB) (Lai and Robbins, 1985). Other popular strategies include Thompson sampling (TS), a Bayesian heuristic introduced by Thompson (1933) that employs randomization to select actions according to the posterior distribution of reward functions. Additionally, the Greedy policy that selects the myopically best action is shown to be effective in contextual bandits (Kannan et al., 2018; Raghavan et al., 2018; Hao et al., 2020; Bastani et al., 2021).
In the linear bandit setting, two types of regret are considered. The Bayesian regret is applicable when the parameter $\theta^{\star}$ is random, and the regret is averaged over three sources of randomness: the observation noise, the randomized algorithm, and the random parameter $\theta^{\star}$. Russo and Van Roy (2014) and Dong and Van Roy (2018) establish an $\widetilde{\mathcal{O}}(d\sqrt{T})$ upper bound for the Bayesian regret of the TS heuristic referred to as LinTS, which matches the minimax optimal bound shown by Dani et al. (2008), where $\widetilde{\mathcal{O}}$ refers to asymptotic order, up to logarithmic factors. On the other hand, the frequentist regret corresponds to scenarios where the underlying parameter $\theta^{\star}$ is fixed, and the expectation is taken only with respect to the randomness of noise and algorithm.
In the frequentist setting, the optimism-based algorithm proposed by Abbasi-Yadkori et al. (2011), known as Optimism in the Face of Uncertainty Linear Bandit Algorithm (OFUL), achieves a $\widetilde{\mathcal{O}}(d\sqrt{T})$ frequentist regret bound, which matches the minimax optimal Bayesian bound. However, for a frequentist variant of LinTS referred to as TS-Freq, Agrawal and Goyal (2013) and Abeille et al. (2017) provide a $\widetilde{\mathcal{O}}(d\sqrt{dT})$ upper bound for the frequentist regret by inflating the posterior distribution. This regret bound falls short of the optimal rate by a factor of $\sqrt{d}$. Recent work by Hamidi and Bayati (2020a) confirms that this inflation is necessary, and thus, the frequentist regret of LinTS cannot be improved. While no general theoretical guarantee exists for the Greedy algorithm in linear bandit problems, it is known that both Greedy and LinTS perform suboptimally in the asymptotic regime (Lattimore and Szepesvari, 2017).
Despite the theoretical gaps in performance, LinTS has shown strong empirical performance, indicating that the inflation of the posterior distribution may be unnecessary in most problem instances, and unfavorable instances are unlikely to occur frequently. Similarly, the Greedy algorithm works well in typical instances (Bietti et al., 2021). Additionally, while OFU-based algorithms are computationally expensive (generally NP-hard as discussed by Dani et al. (2008); Russo and Van Roy (2014); Agrawal (2019)), LinTS and Greedy are known for their computational efficiency. This unsettling disparity between theoretical, computational, and empirical performance has motivated our investigation into the following two questions: Is it possible to identify problematic instances where LinTS and Greedy could potentially fail, and apply a “course-correction” to ensure competitive frequentist regret bounds? And can this be achieved without compromising their impressive empirical performance and computational efficiency? In this paper, we provide positive answers to both questions. Specifically, we make the following contributions.
  • Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid. This technique is essential in maximizing the utilization of historical data, thereby enabling the “course-correction”.
  • We derive a general instance-dependent frequentist regret bound for the proposed POFUL algorithms, a family of algorithms including OFUL, LinTS, TS-Freq, and Greedy as special cases. The bound can be efficiently calculated in a data-driven manner based on the observation history.
  • We propose variants of LinTS and Greedy that enjoy the minimax optimal frequentist regret, while retaining most of the desirable properties such as computational efficiency.
The distinguishing feature of our study from the existing literature is the geometric analysis of the $d$-dimensional confidence ellipsoid. This is in stark contrast to techniques outlined in previous studies that may inadvertently discard valuable information as a consequence of their overly narrow concentration on comparatively rudimentary summary statistics.
Related literature.
The ideas presented in this paper bear similarities to existing literature that explores different forms of exploration (Russo and Van Roy, 2016; Kirschner and Krause, 2018) or recent works by (Bastani et al., 2021; Pacchiano et al., 2020; Hamidi and Bayati, 2020b) that aim to utilize observed data for reducing exploration. However, the methodology and the manner in which the data is utilized in this paper differ significantly. For instance, Bastani et al. (2021) examines the minimum eigenvalue of the sample covariance matrix which is a single-parameter summary statistic of the observed data, while Hamidi and Bayati (2020b) incorporates information from the (one-dimensional) confidence intervals around the reward values. In contrast, our approach leverages the complete geometric information of the $d$-dimensional confidence ellipsoid.
2 Setup and Preliminaries
Notations.
[…]
Problem Formulation.
We consider the stochastic linear bandit problem. Let $\theta^{\star}\in\mathbb{R}^{d}$ be a fixed but unknown parameter. At each time $t\in[T]$, a policy $\pi$ selects action $x_{t}$ from a set of action $\mathcal{X}_{t}\subset\mathbb{R}^{d}$ according to the past observations and receives a reward $r_{t}=\langle{x_{t}},{\theta^{\star}}\rangle+\varepsilon_{t}$, where $\varepsilon_{t}$ is mean-zero noise with a distribution specified in Assumption 3 below. We measure the performance of $\pi$ with the cumulative expected regret $\mathcal{R}(T)=\sum_{t=1}^{T}\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x_{t}},{\theta^{\star}}\rangle,$ where $x^{\star}_{t}$ is the best action at time $t$, i.e., $x^{\star}_{t}=\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$. Let $\mathcal{F}_{t}$ be a $\sigma$-algebra generated by the history $(x_{1},r_{1},\ldots,x_{t},r_{t})$ and the prior knowledge, $\mathcal{F}_{0}$. Therefore, $\{\mathcal{F}_{t}\}_{t\geq 0}$ forms a filteration such that each $\mathcal{F}_{t}$ encodes all the information up to the end of period $t$.
2.1 Assumptions
We make the following assumptions that are standard in the relevant literature.
Assumption 1 (Bounded parameter)
The unknown parameter $\theta^{\star}$ is bounded as $\|\theta^{\star}\|\leq S$, where $S>0$ is known.
Assumption 2 (Bounded action sets)
The action sets $\{\mathcal{X}_{t}\}$ are uniformly bounded and closed subsets of $\mathbb{R}^{d}$, such that $\|x\|\leq X_{t}$ for all $x\in\mathcal{X}_{t}$ and all $t\in[T]$, where $X_{t}$’s are known and $\sup_{t\geq 1}\left\{X_{t}\right\}<\infty$. Without loss of generality, we normalize the actions such that $\mathcal{X}_{t}\subset{\mathcal{S}}_{d-1}$ for all $t\in[T]$.
Assumption 3 (Noise)
The noise sequence $\{\varepsilon_{t}\}_{t\geq 1}$ is conditionally mean-zero and $R$-subgaussian, where $R$ is known. Formally, for all real valued $\lambda$, $\mathbb{E}\left[e^{\lambda\varepsilon_{t}}|\mathcal{F}_{t}\right]\leq\exp\left(\lambda^{2}R^{2}/2\right)$. This condition implies that $\mathbb{E}\left[\varepsilon_{t}|\mathcal{F}_{t}\right]=0$ for all $t\geq 1$.
2.2 Regularized Least Square and Confidence Ellipsoid
Here, we review useful frequentist tools developed in Abbasi-Yadkori et al. (2011).
Let $(x_{1},\dots,x_{t})$ be an arbitrary sequence of actions and $(r_{1},\dots,r_{t})$ be the corresponding rewards. In LB problems, the parameter $\theta^{*}$ is estimated with the regularized ordinary least-square estimator. Fix a regularization parameter $\lambda$, the design matrix, and the Regularized Least Square (RLS) estimate are respectively given by
$V_{t}=\lambda I+\sum_{s=1}^{t}x_{s}x_{s}^{\top},\quad\widehat{\theta}_{t}=V_{t}^{-1}\sum_{s=1}^{t}x_{s}r_{s}.$ (1)
The following proposition in Abbasi-Yadkori et al. (2011) states that the RLS estimate $\widehat{\theta}_{t}$ concentrates around $\theta^{\star}$ properly.
Proposition 1 (Theorem 2 in Abbasi-Yadkori et al. (2011))
Fix $\delta\in(0,1)$, it holds with probability $1-\delta$ that
$\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta),\quad\left|\langle{x},{\widehat{\theta}_{t}-\theta^{\star}}\rangle\right|\leq\|x\|_{V_{t}^{-1}}\beta_{t}^{RLS}(\delta)$
for all $x\in\mathbb{R}^{d}$, where
$\beta_{t}^{RLS}(\delta)=R\sqrt{2\log\frac{(\lambda+t)^{d/2}\lambda^{-d/2}}{\delta}}+\sqrt{\lambda}S.$ (2)
Hereafter, we might omit the dependence of $\beta_{t}$ on $\delta$ if there is no ambiguity. Propostion 1 enables us to construct the following sequence of confidence ellipsoids.
Definition 1
Fix $\delta\in(0,1)$. We define the RLS confidence ellipsoid as $\mathcal{E}_{t}^{RLS}(\delta)=\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta)\}$.
The next proposition, known as the elliptical potential lemma, presents a concentration result that plays a central role in bounding the regret. Basically, this proposition is saying that the cumulated prediction error incurred by the action sequence used to estimate $\theta^{\star}$ is small.
Proposition 2 (Lemma 11 in Abbasi-Yadkori et al. (2011))
If $\lambda>1$, for an arbitrary sequence $(x_{1},\dots,x_{t})$, it holds that $\sum_{s=1}^{t}\left\|x_{s}\right\|_{V_{s}^{-1}}^{2}\leq 2\log\frac{\operatorname{det}\left(V_{t+1}\right)}{\operatorname{det}(\lambda I)}\leq 2d\log\left(1+\frac{t}{\lambda}\right).$
3 POFUL Algorithms
In this section, we introduce a generalized framework of OFUL that we call POFUL (Pivot OFUL). POFUL takes as input a sequence of inflation parameters $\{\iota_{t}\}_{t\in[T]}$, feasible (randomized) pivots $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ and optimism parameters $\{\tau_{t}\}_{t\in[T]}$. The inflation parameters are used to construct confidence ellipsoids that contain $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ with high probability. This turns into the following condition for $\{\iota_{t}\}_{t\in[T]}$ and $\{\widetilde{\theta}_{t}\}_{t\in[T]}$.
Definition 2
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. Given the inflation parameters $\{\iota_{t}\}_{t\in[T]}$, we call random variables $\{\widetilde{\theta}_{t}\}_{t\in[T]}$ feasible pivots if for all $t\in[T]$,
$\mathbb{P}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]\geq 1-\delta^{\prime},$
where we define the pivot ellipsoid as $\mathcal{E}_{t}^{PVT}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\iota_{t}\beta^{RLS}_{t}(\delta)\right\}$.
At each time $t$, POFUL chooses the action that maximizes the optimistic reward
$\widetilde{x}_{t}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\ \langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$ (3)
A pseudocode representation of POFUL is provided in Algorithm 1.
Algorithm 1 POFUL
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\{\tau_{t}\}_{t\in[T]}$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Sample a feasible pivot $\widetilde{\theta}_{t}$ with respect to $\iota_{t}$ according to Definition 2
     $\widetilde{x}_{t}\leftarrow\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+\widetilde{x}_{t}\widetilde{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}\widetilde{x}_{s}r_{s}$.
  end for
The intuition behind POFUL is combining the exploration mechanism of OFUL and LinTS. OFUL encourages exploration by introducing the uncertainty term $\tau_{t}\|x\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})$ in the reward, while LinTS explores by sampling from the confidence ellipsoid. We allow POFUL to select an arbitrary pivot from $\mathcal{E}_{t}^{PVT}$ to encompass arbitrary exploration mechanisms within $\mathcal{E}_{t}^{PVT}$. By definition, POFUL includes OFUL, LinTS, TS-Freq, and Greedy as special cases.
OFUL
For stochastic linear bandit problems, OFUL chooses actions by solving the optimization problem $\max_{x\in\mathcal{X}_{t}}\ \langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$ Therefore, OFUL is a specially case of POFUL where $\iota_{t}=0$, $\tau_{t}=1$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$, the center of the confidence ellipsoid, for all $t\in[T]$.
LinTS and TS-Freq
Linear Thompson Sampling (LinTS) algorithm is a generic randomized algorithm that samples from a distribution constructed from the RLS estimate at each step. At time $t$, LinTS samples as $\widetilde{\theta}_{t}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$, where $\delta^{\prime}=\delta/2T$, and $\iota^{TS}_{t}$ is an inflation parameter controlling the scale of the sampling range, and $\eta_{t}$ is a random sample from a normalized sampling distribution $\mathcal{D}^{SA}(\delta^{\prime})$ that concentrates with high probability.
Definition 3
Fix any $\delta\in(0,1)$. $\mathcal{D}^{SA}(\delta)$ is a distribution such that $\mathbb{P}_{\theta\sim\mathcal{D}^{SA}(\delta)}\left[\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq 1\right]\geq 1-\delta.$
LinTS is a special case of POFUL where $\iota_{t}=\iota^{TS}_{t}$, $\tau_{t}=0$ and $\widetilde{\theta_{t}}=\widehat{\theta}_{t}+\iota^{TS}_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$. We remark that setting the inflation parameter $\iota_{t}=\widetilde{\mathcal{O}}(1)$ corresponds to the original LinTS algorithm. On the other hand, setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$ corresponds to the frequentist variant of LinTS studied in Agrawal and Goyal (2013); Abeille et al. (2017), denoted by TS-Freq. This means TS-Freq inflates the posterior by a factor of order $\sqrt{d}$.
Greedy
Greedy is a special case of POFUL where $\iota_{t}=\tau_{t}=0$ and $\widetilde{\theta}_{t}=\widehat{\theta}_{t}$ for all $t\in[T]$.
4 Frequentist Regret Analysis of POFUL
In this section, we present the frequentist regret analysis of POFUL algorithms. We first introduce useful concentration events that hold with high-probability.
Definition 4
Fix $\delta\in(0,1)$ and $\delta^{\prime}=\delta/2T$. We define $\beta^{PVT}_{t}(\delta^{\prime})\coloneqq\iota_{t}(\delta)\beta^{RLS}_{t}(\delta^{\prime})$ and
$\ \widehat{\mathcal{A}}_{t}$$\coloneqq\left\{\forall s\leq t:\left\|\widehat{\theta}_{t}-\theta^{\star}\right\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\},\quad\widetilde{\mathcal{A}}_{t}\coloneqq\left\{\forall s\leq t:\left\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\right\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})\right\}.$
We also define $\mathcal{A}_{t}\coloneqq\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$
Proposition 3
Under Assumptions 1, 2 and 3 , we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta.$
In the following, we decompose the regret of POFUL into three terms, $\mathcal{R}^{PE}(T)$, $\mathcal{R}^{PVT}(T)$ and $\mathcal{R}^{RLS}(T)$, and sketch how they are bounded separately. Let $x_{t}^{\star}=\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta^{\star}}\rangle$ be the benchmark action and $\widetilde{x}_{t}$ be the action chosen by POFUL, the regret decomposes as
$\mathcal{R}(T)$ $=\sum_{t=1}^{T}\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle+\sum_{t=1}^{T}\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$
$\eqqcolon\mathcal{R}^{PE}(T)+\mathcal{R}^{PVT}(T)+\mathcal{R}^{RLS}(T).$
$\mathcal{R}^{RLS}(T)$ is the regret due to the estimation error of the RLS estimator. Note that $\widetilde{x}_{t}$’s are the action sequence used to construct the RLS estimate. By Proposition 2 their cumulative $V_{t}^{-1}$-norm is bounded by $2d\log(1+t/\lambda)$. Hence the upper bound of $\mathcal{R}^{RLS}(T)$ is essentially determined by the $V_{t}$-distance between $\widehat{\theta}_{t}$ and $\theta^{\star}$, which are characterized by the confidence ellipsoids $\mathcal{E}^{RLS}_{t}$. $\mathcal{R}^{PVT}(T)$ corresponds to the regret due to the exploration of POFUL by choosing the pivot parameter $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$. Similar to $\mathcal{R}^{RLS}(T)$, the upperbound is related to the $V_{t}$-distance between $\widetilde{\theta}_{t}$ and $\widehat{\theta}_{t}$, which is controlled by construction and depends on the inflation parameter $\iota_{t}$. As a result, it can be shown that $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$ and $\mathcal{R}^{PVT}(T)=\widetilde{\mathcal{O}}(d(\sum_{t=1}^{T}\iota_{t}^{2})^{1/2})$ with probability $1-\delta$. Notably, $\mathcal{R}^{RLS}(T)=\widetilde{\mathcal{O}}(d\sqrt{T})$, matching the known minimax lowerbound. For $\mathcal{R}^{PVT}(T)$ to match this lower bound as well, one way is to set $\iota_{t}=\widetilde{\mathcal{O}}(1)$ for all $t\in[T]$. However, later we see this could be problematic when bounding $\mathcal{R}^{PE}(T)$.
$\mathcal{R}^{PE}(T)$ corresponds to the pessimism term. In the Bayesian analysis of LinTS, the expectation of this term is 00, with respect to $\theta^{\star}$, as long as $\widetilde{\theta}_{t}$ and $\theta^{\star}$ are sampled from the same distribution which occurs when LinTS has access to the true prior distribution for $\theta^{\star}$. In the frequentist analysis, however, we need to control the pessimism term incurred by any random sample $\widetilde{\theta}_{t}$. For OFUL, this term is bounded properly with high probability by the construction of OFUL actions. For LinTS, the only known analysis due to Agrawal and Goyal (2013); Abeille et al. (2017) gives a bound of order $\widetilde{\mathcal{O}}(d\sqrt{dT})$ using an optimism-based approach, which is worse than the Bayesian regret by a factor of $\sqrt{d}$. The key component of their proof is introducing inflation to the posterior variance by setting $\iota_{t}=\widetilde{\mathcal{O}}(\sqrt{d})$. In Abeille et al. (2017), the inflation is necessary for them to lower bound the probability that LinTS samples optimistic $\widetilde{\theta}_{t}$ that leads to a nonpositive pessimism i.e. $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle$. For Greedy, there isn’t a general theoretical guarantee to bound the pessimism term without additional structures. The way of bounding it distinguishes our method from the methods based on optimism (Agrawal and Goyal, 2013; Abeille et al., 2017), and enables us to improve the frequentist regret of LinTS and Greedy.
In fact, non-optimistic samples might be able to control the pessimism term. Our method does not force choosing optimistic samples and hence there is no need to inflate the posterior. As a result, it does not introduce a $\sqrt{d}$-gap in the regret. Instead, we introduce a measurement of the “quality” of POFUL actions with respect to the regret, and develop a computable upper bound for this quality using geometric approaches. The main merits are an instance-dependent regret bound for POFUL that matches the minimax optimal regret in some settings, and variants of LinTS and Greedy with provable frequentist regret bound that is minimax optimal up to only a constant factor.
4.1 An Instance-Dependent Regret Bound for POFUL
In the following, we condition on the event $\mathcal{A}_{T}$ which holds with probability $1-\delta$. The following proposition bounds the instantaneous regret of POFUL and we defer the proof to Appendix D.2.
Proposition 4
Suppose ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
On the RHS, since the oracle optimal action sequence $\{x^{\star}_{t}\}_{t\in[T]}$ is unknown to the algorithm and is different from the action sequence $\{\widetilde{x}_{t}\}_{t\in[T]}$ played by POFUL, one cannot apply Proposition 2 to bound the summation $\sum_{t=1}^{T}\left\|\widetilde{x}_{t}\right\|_{V_{t}^{-1}}^{2}$ to get an upperbound of the regret. To fix the problem due to this discrepancy, the key point is connecting $\{\widetilde{x}_{t}\}_{t\in[T]}$ and $\{x^{\star}_{t}\}_{t\in[T]}$ in terms of the $V_{t}^{-1}$ norm.
Definition 5
For each $t\geq 1$, let $\widetilde{x}_{t}$ and $x_{t}^{\star}$ respectively denote the action chosen by POFUL and the optimal action. We define the sensitivity ratio at time $t$ as $\alpha_{t}\coloneqq\|x_{t}^{\star}\|_{V_{t}^{-1}}/\|\widetilde{x}_{t}\|_{V_{t}^{-1}}$. We also define the regret ratio at time $t$ as $\mu_{t}\coloneqq\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$.
Recall the Cauchy–Schwarz inequality $\langle{x},{\theta_{1}-\theta_{2}}\rangle\leq\|x\|_{V_{t}^{-1}}\|\theta_{1}-\theta_{2}\|_{V_{t}}$ , intuitively $\alpha_{t}$ serves as a measure of how sensitive the reward $\langle{\widetilde{x}_{t}},{\theta}\rangle$ is compared with that of the optimal action $\langle{x^{\star}_{t}},{\theta}\rangle$, with respect to the change of the parameter $\theta$ in the $V_{t}$-norm.
The intuition behind the definition for $\mu_{t}$ is connecting POFUL and OFUL. It provides an upper bound for the ratio of instantaneous regrets incurred by POFUL and the OFUL-type upper bound. Specifically, Proposition 4 indicates $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq\mu_{t}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$, and we can check that the instantaneous regret of OFUL satisfies $\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq 2\|\widetilde{x}_{t}\|_{V_{t}^{-1}}\beta^{RLS}_{t}$. In this sense, our definition of $\mu_{t}$ and $\alpha_{t}$ can be extended to OFUL. One feasible choice is setting $\mu_{t}\equiv 2$ for OFUL and therefore $\alpha_{t}=(\mu_{t}-1-\iota_{t}-\tau_{t})/(1+\iota_{t}-\tau_{t})=1$ (recall $\iota_{t}=0$ and $\tau_{t}=1$ for OFUL). The following Theorem connects $\{\mu_{t}\}_{t\in[T]}$ and $\mathcal{R}(T)$ and we postpone the proof to Appendix D.3.
Theorem 1 (Oracle frequentist regret bound for POFUL)
Fix $\delta\in(0,1)$ and let $\delta^{\prime}=\delta/2T$. Under Assumptions 1, 2 and 3, with probability $1-\delta$, POFUL achieves a frequentist regret of
$\mathcal{R}(T)\leq\sqrt{2d\left(\sum_{t=1}^{T}\mu_{t}^{2}\right)\log\left(1+\frac{T}{\lambda}\right)}\beta_{T}^{RLS}(\delta^{\prime}).$ (4)
Remark 1
A similar notion of (the reciprocal of) our $\alpha_{t}$ can be found in the Discussion section in Abeille et al. (2017), where they indicate their proof of regret bound can be reproduced if the algorithm samples $\widetilde{\theta}_{t}$ that satisfies $\|x^{\star}(\widetilde{\theta}_{t})\|_{V_{t}^{-1}}\geq\alpha\|x^{\star}(\theta^{\star}_{t})\|_{V_{t}^{-1}}$ with constant probability for some $\alpha>0$. They comment that the set of these parameters doesn’t coincide with the set of optimistic parameters and it remains an open problem whether one could relax the requirement of oversampling the posterior through this observation. Our work answers this question positively in the next section by investigating the property of (the reciprocal of) $\alpha$ using geometric arguments, providing an explanation for LinTS’s empirical success.
Remark 2
We call Theorem 1 an oracle regret bound as $\{\mu_{t}\}_{t\in[T]}$ for general POFUL depends on the unknown system parameter $\theta^{\star}$. In general, they cannot be calculated by the decision-maker. Nevertheless, note that $(1+\iota_{t}-\tau_{t})\geq 0$ for OFUL, LinTS, and Greedy. Suppose we can derive computable upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$ for $\{\alpha_{t}\}_{t\in[T]}$ respectively, then we could calculate upper bounds for $\{\mu_{t}\}_{t\in[T]}$ as well, and Theorem 1 instantly turns into a data-driven regret bound for POFUL. This is indeed our aim in the next section.
5 A Data-Driven Approach to Calibrating POFUL
In this section, we present the main contribution of this work which provides a data-driven approach to calibrating POFUL. We’ll focus on LinTS and Greedy as they lack theoretical results of frequentist regret bound. In these settings, we have $\tau_{t}=0$ and hence $\mu_{t}\coloneqq(1+\alpha_{t})(1+\iota_{t})$. Note that $\iota_{t}$ is the inflation parameter that can be controlled by a decision-maker, the essential point is to find a computable, non-trivial upper bound for the sensitivity ratio $\alpha_{t}$, which turns into an upper bound for the regret ratio $\mu_{t}$ that’s deeply related to the frequentist regret of POFUL.
In the following, we first construct such upper bounds $\{\widehat{\alpha}_{t}\}_{t\in[T]}$. Our strategy capitalizes on geometric insights related to the properties of the confidence ellipsoids. Subsequently, we show how these bounds can be utilized to calibrate POFUL. Specifically, we propose the TS-MR and Greedy-MR algorithms that attain a provable minimax-optimal frequentist regret up to constant factors, and yet keep the majority of their original properties such as computational efficiency and low regret in most instances.
5.1 Upper Bounding the Sensitivity Ratio
We consider two scenarios: (1) Continuous action space and (2) Discrete action space. For sake of a better illustration, we consider $\mathcal{X}_{t}\equiv{\mathcal{S}}_{d-1}$ for the first scenario, which is a standard example of continuous action space, and is the same as the setting considered in Abeille et al. (2017). The analysis of the second scenario is deferred to Appendix A.
Continuous action space.
Note that in this case, the optimal action $x_{t}^{\star}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$ takes a closed form $x^{\star}_{t}(\theta)=\theta/\|\theta\|$, we only need to study the property of the confidence ellipsoids of $\theta$, namely, $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$. A direct idea is to consider respectively the smallest and largest value of $\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}$ for $\theta$ in the ellipsoids. Specifically, we have
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}.$ (5)
At first glance, one might think that (5) is too conservative to provide useful calibration for POFUL. However, it turns out that the lower bound is pretty tight in practice. As is illustrated in Figure 1. The intuition is that, when POFUL has implemented sufficient exploration so that the confidence ellipsoid $\mathcal{E}_{t}$ is small enough, the potentially optimal actions concentrate in a small cap $\mathcal{C}_{t}$ on the unit hypersphere ${\mathcal{S}}_{d-1}$. All actions within $\mathcal{C}_{t}$ roughly point to the same direction and thus have similar $V_{t}^{-1}$-norm. The following main theorem of this paper derives a upper bound of $\alpha_{t}$ based on this idea. We sketch the proof in Appendix B.
Figure 1: Illustration of the set $\mathcal{C}_{t}$ of potentially optimal actions in $\mathbb{R}^{2}$.
The following assumption simplifies the proof, and we remark that it holds almost surely.
Assumption 4
For all $t$, $V_{t}$ has distinct eigenvalues, i.e. $\lambda_{1}(V_{t})>\lambda_{2}(V_{t})\dots>\lambda_{d}(V_{t})\geq 0$.
Theorem 2
Under Assumpsion 4, define $\beta^{PVT}_{t}=\iota_{t}\beta^{RLS}_{t}$ and
[…]
where we let
$\quad m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t}/\lambda_{d}(V_{t}))^{2},\quad M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t})^{2})$
and $k\in[d]$ is such that $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$. Then for all $t\in[T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds for all $s\leq t$ that
$\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s},\quad\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s},\quad\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}.$
To better understand what Theorem 2 implies, we discuss some special cases.
Case 1: a pure exploration regime.
When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$. Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$. Specifically, if $V_{t}=D\mathbb{I}_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
Case 2: linear structures.
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$. To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding sensitivity ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
Proposition 5
Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{\max}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
5.2 TS-MR and Greedy-MR Algorithms
In the following, we propose the algorithm Linear Thompson Sampling with Maximum (Regret) Ratio (TS-MR). The intuition is to detect the performance of LinTS by estimating $\mu_{t}$ and avoid bad LinTS actions by switching to OFUL occasionally.
Specifically, at each time $t$, TS-MR calculates an estimate $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Appendix E.
By construction, TS-MR and Greedy-MR have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
Corollary 1
TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
6 Simulations
(a) example 1
(b) example 2
(c) example 3
Figure 2: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
Example 1. Stochastic linear bandit with uniformly and independently distributed actions.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm. At each time $t$, we generate 100 i.i.d. random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$. This is a basic example of standard stochastic linear bandit problems without any extra structure. We set the threshold $\mu=8$ for TS-MR and Greedy-MR. We average simulation results over 20 independent runs. The results are summarized in Figure 1(a). TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
Example 2. Contextual bandits embedded in the linear bandit problem (Abbasi-Yadkori, 2013).
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$. At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0. Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks. This problem is equivalent to a 10-armed contextual bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. Again, we average simulation results over 20 independent runs. The results are summarized in Figure 1(b). In this setting, Greedy performs suboptimally due to a lack of exploration for all arms. Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects bad regret ratio $\widehat{\mu}_{t}$.
Example 3. Prior mean mismatch (Hamidi and Bayati, 2020a).
This is an example in which LinTS is shown to incur linear Bayesian regret. We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$. It is shown in Hamidi and Bayati (2020a) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$. Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action. We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. We carried out 20 independent runs. The results are summarized in Figure 1(c). We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
7 Conclusion
In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, and Greedy as special cases. Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm. Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest. As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail. We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
Appendix A Discrete Action Space
In this scenario, we consider upper bounding the sensitivity ratio for discrete action space. Compared with the unit hypersphere ${\mathcal{S}}_{d-1}$ considered in Section 5.1 , the main difference is that the oracle optimal action $\theta/\|\theta\|$ for $\theta$ might not lie in the action space $\mathcal{X}_{t}$, so we don’t have a closed form for the attainable optimal action $x^{\star}_{t}(\theta)\coloneqq\mathop{\mathrm{argmax}}_{x\in\mathcal{X}_{t}}\langle{x},{\theta}\rangle$.
To tackle this problem, we need to characterize the property of all potentially optimal actions. We provide two approaches to solve this problem. This first approach works for the setting where the cardinalities of the action sets are relatively small. The second approach utilizes the statistical properties of the action sets and works for the setting where the cardinalities of the action sets are large.
Small cardinality setting.
Note that the optimistic reward of POFUL when $\tau_{t}=0$ is $\langle{x},{\widetilde{\theta}_{t}}\rangle$, where $\widetilde{\theta}_{t}\in\mathcal{E}^{PVT}_{t}$ is an arbitrary pivot parameter. We have $L_{t}(x)\leq\langle{x},{\widetilde{\theta}_{t}}\rangle\leq U_{t}(x)$, where we define the confidence lower and upper bounds as
$L_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle-\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}),\quad U_{t}(x)=\langle{x},{\widehat{\theta}_{t}}\rangle+\iota_{t}\|x\|_{V_{t}^{-1}}\beta^{RLS}_{t}(\delta^{\prime}).$
The potentially optimal $x$ that might optimize $\langle{x},{\widetilde{\theta}_{t}}\rangle$ should satisfy
$U_{t}(x)\geq\max_{y\in\mathcal{X}_{t}}\ L_{t}(y),$
i.e. $x$ is not dominated by any other action in $\mathcal{X}_{t}$. Therefore, define
$\mathcal{C}_{t}(\beta)\coloneqq\{x\in\mathcal{X}_{t}:\langle{x},{\widehat{\theta}_{t}}\rangle+\|x\|_{V_{t}^{-1}}\beta\geq\max_{y\in\mathcal{X}_{t}}\ \langle{y},{\widehat{\theta}_{t}}\rangle-\|y\|_{V_{t}^{-1}}\beta\}.$
An upper bound for $\alpha_{t}$ can be calculated as
$\alpha_{t}\leq\widehat{\alpha}_{t}\coloneqq\frac{\sup_{x\in\mathcal{C}_{t}(\beta_{t}^{RLS}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}{\inf_{x\in\mathcal{C}_{t}(\beta_{t}^{PVT}(\delta^{\prime}))}\|x\|_{V_{t}^{-1}}}.$ (6)
Large cardinality setting.
In this setting, we characterize the distance between $\theta/\|\theta\|$ and $x^{\star}_{t}(\theta)$. We first introduce the notion of Euclidean covering radius.
Definition 6
The Euclidean covering radius $\rho_{t}$ of a set $\mathcal{X}_{t}\subseteq{\mathcal{S}}_{d-1}$ is given by
$\rho_{t}=\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\coloneqq\max_{y\in{\mathcal{S}}_{d-1}}\min_{x\in\mathcal{X}_{t}}\|y-x\|.$
With the help of the Euclidean covering radius, we prove the following variant of Theorem 2. The proof is postponed to Appendix D
Theorem 3
Under Assumpsion 4, let $\{\rho_{t}\}$ be the sequence of Euclidean covering radius of $\mathcal{X}_{t}$, define $\Psi_{t},\Phi_{t}$ as in Theorem 2. Then for all $t\in[1,T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$ , it holds that
$\alpha_{s}\leq\frac{\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}{\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})}$ (7)
for all $s\leq t$.
To justify this new lower bound, it remains to show that $\rho_{t}$ is small enough under proper conditions. For the sake of a better illustration, we focus on action sets that consist of uniformly and independently distributed points on ${\mathcal{S}}_{d-1}$. The following proposition shows that the Euclidean covering radius has an high-probability upper bound that decays at the order of $\widetilde{\mathcal{O}}(N^{-1/d+o(1)})$, where $N$ is the cardinality of actions.
Proposition 6 (Corollary 2.9 in Reznikov and Saff [2016])
Suppose $\mathcal{X}$ is a set of $N$ uniformly and independently distributed random points on the unit sphere ${\mathcal{S}}_{d-1}$, for any $\kappa>0$, there exist constants $C_{\kappa,d}>0$, such that
$\mathbb{P}\left[\rho(\mathcal{X}_{t},{\mathcal{S}}_{d-1})\leq C_{\kappa,d}\left(\frac{\log N}{N}\right)^{1/d}\right]\geq 1-N^{-\kappa}.$
Appendix B Sketch of the Proof of Theorem 2
In this section, we present the proof of Theorem 2. In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
The proof of Theorem 2 decomposes into two steps. First, we cut out a special hypersurface $\mathcal{H}_{t}$ within $\mathcal{E}_{t}$ and show that for all $\theta\in\mathcal{H}_{t}$, the corresponding optimal action $x^{\star}(\theta)$ has $V_{t}$-norm bounded from above and below. Note that the set of optimal actions for $\theta\in\mathcal{H}_{t}$ coincides with that for $\theta\in\mathcal{E}_{t}$, we get upper and lower bounds for $V_{t}$-norm of all potential actions in the ellipsoid. Second, we show that upper and lower bounds for the $V_{t}$-norm can be transferred into upper and lower bounds for the $V_{t}^{-1}$-norm. Hence, we get an upper bound for $\alpha_{t}$ by calculating the ratio of the upper bound to the lower bound. We postpone the detailed proof for all lemmas in this section to Appendix C.
Step 1. Bounding the $V_{t}$-norm of actions.
We illustrate our geometric approach in $\mathbb{R}^{2}$ in Figure 3. We first upper bound the $V_{t}$ norm for optimal actions corresponding to any $\theta\in\mathcal{E}_{t}$. In the following, we define $\phi_{t}=\sqrt{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}-\beta_{t}^{2}}$.
Lemma 1
For any $\theta\in\mathcal{E}_{t}$, we have $\|x^{\star}(\theta)\|_{V_{t}}\leq\|\widehat{\theta}_{t}\|_{V_{t}^{2}}/\phi_{t}$.
The proof is based on the observation that for any $\theta\in\mathcal{E}_{t}$, the line $\{t\theta:t\in\mathbb{R}\}$ (orange dashed lines in Figure 3) and ellipsoid $\mathcal{E}_{t}$ (in green) must have at least one intersection point.
In order to lower bound the $V_{t}$-norm, we define the hypersurface $\mathcal{H}_{t}\coloneqq\mathcal{E}_{t}\cap\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$, i.e., the intersection of the interior of the confidence ellipsoid $\mathcal{E}_{t}$ and the ellipsoid $\left\{\theta:\|\theta\|_{V_{t}}=\phi_{t}\right\}$ (blue dashed curve). We can check this intersection is non-empty since $\phi_{t}\widehat{\theta}_{t}/\|\widehat{\theta}_{t}\|_{V_{t}}\in\mathcal{E}_{t}$. A lower bound for the $V_{t}$-norm is given by the following lemma.
Lemma 2
For any $\theta\in\mathcal{H}_{t}$, we have
$\|x^{\star}(\theta)\|_{V_{t}}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
The proof is based on the fact that for any $\theta\in\mathcal{H}_{t}$, we have $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$ and $\|\theta\|_{V_{t}}=\phi_{t}$. Also recall $x^{\star}(\theta)=\theta/\|\theta\|$ and hence $\|x^{\star}(\theta)\|_{V_{t}}=\|\theta\|_{V_{t}}/\|\theta\|$.
Figure 3: Illustration of Step 1 in $\mathbb{R}^{2}$.
Step 2. Bounding the $V_{t}^{-1}$-norm of actions
The following lemma bounds the $V_{t}^{-1}$-norm of an action $x$ given a range of its $V_{t}$-norm.
Lemma 3
Suppose $V$ has distinct eigenvalues $\lambda_{1}(V)>\lambda_{2}(V)\dots>\lambda_{d}(V)$ and $x\in\mathcal{B}_{d}$ satisfies $m\leq\|x\|_{V}^{2}\leq M$, where $M\geq 0$ is such that $\lambda_{k}(V)\leq M<\lambda_{k+1}(V)$. We have
$\frac{1}{\lambda_{k}}+\frac{1}{\lambda_{k+1}}-\frac{M}{\lambda_{k}\lambda_{k+1}}\leq\|x\|_{V^{-1}}^{2}\leq\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$ (8)
The proof is based on writing $x$ as a linear combination of the eigenvectors of $V_{t}$, $x=\sum_{i=1}^{d}w_{i}u_{i}$, and convert the original problem into a linear programming problem with constraints using the expressions $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
Combining the two steps, we complete the proof of Theorem 2.
Appendix C Proof of Lemmas for Theorem 2
C.1 Proof of Lemma 1
Let $\theta=tx$ where $x$ is any unit vector in $\mathbb{R}^{d}$ and $t\in\mathbb{R}$ is a scalar. Consider the equation that characterizes the intersection $\{tx:t\in\mathbb{R}\}\cap\mathcal{E}_{t}$, namely $(tx-\widehat{\theta}_{t})^{\top}V_{t}(tx-\widehat{\theta}_{t})=\beta_{t}^{2}$. Equivalently, we have $t^{2}\|x\|_{V_{t}}^{2}-2tx^{\top}V_{t}\widehat{\theta}_{t}+\phi_{t}^{2}=0$. This quadratic equation of $t$ has at least one root if the discriminant is non-negative, i.e. $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$. Then by direct computation,
$\|x\|_{V_{t}}\leq\frac{\sqrt{(x^{\top}V_{t}\widehat{\theta}_{t})^{2}}}{\phi_{t}}\leq\frac{\sqrt{x^{\top}x}\sqrt{(\widehat{\theta}_{t})^{\top}V_{t}^{\top}V_{t}\widehat{\theta}_{t}}}{\phi_{t}}=\frac{\|\widehat{\theta}_{t}\|_{V_{t}^{2}}}{\phi_{t}}.$
Note that $\theta\in\mathcal{E}_{t}$ if and only if $\theta=tx$ for some $x$ that satisfies $4(x^{\top}V_{t}\widehat{\theta})^{2}\geq 4\|x\|_{V_{t}}^{2}\phi_{t}^{2}$, we conclude the proof. $\square$
C.2 Proof of Lemma 2
Note that for any $\theta\in\mathcal{H}_{t}\subset\mathcal{E}_{t}$, it holds that $\|\theta\|\leq\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})$. Also, by the construction of $\mathcal{H}_{t}$, we have $\|\theta\|_{V_{t}}=\phi_{t}$. Then we conclude the proof by the definition
$\|x^{\star}(\theta)\|_{V_{t}}=\frac{\|\theta\|_{V_{t}}}{\|\theta\|}\geq\frac{\phi_{t}}{\|\widehat{\theta}_{t}\|+\beta_{t}/\lambda_{d}(V_{t})}.$
$\square$
C.3 Proof of Lemma 3
Let $V=U\Lambda U^{\top}$ be the eigendecomposition of $V$, so that the columns of $U$, $\{u_{1},\dots,u_{d}\}$, form an orthonormal basis of $\mathbb{R}^{d}$ and $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$ is a diagonal matrix. We can write $x$ as a linear combination of $\Lambda=\text{diag}(\lambda_{1},\dots,\lambda_{d})$: $x=\sum_{i=1}^{d}w_{i}u_{i}$, where $\sum_{i=1}^{d}w_{i}^{2}=1$. By direct computation, we have $\|x\|_{V}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}$ and $\|x\|_{V^{-1}}^{2}=\sum_{i=1}^{d}w_{i}^{2}\lambda_{i}^{-1}$.
We would like to study the range of $\|x\|_{V^{-1}}^{2}$ when $\|x\|_{V}^{2}$ is bounded as $m\leq\|x\|_{V}^{2}\leq M$. First, let’s focus on maximizing $\|x\|_{V^{-1}}^{2}$, it suffices to solve the LP problem
maximize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}\quad\text{s.t.}\quad\forall i,a_{i}\geq 0,\ \sum_{i=1}^{d}a_{i}=1,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\ \sum_{i=1}^{d}a_{i}\lambda_{i}\leq M.$
The Lagrangian is given by
$L(d,\mu,\eta)=\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}+\mu(1-\sum_{i=1}^{d}a_{i})+\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)+\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})+\sum_{i=1}^{d}\kappa_{i}a_{i}.$
The KKT conditions are given by
$\begin{cases}&\nabla_{a_{i}}L=\lambda_{i}^{-1}-\mu+\eta\lambda_{i}-\gamma\lambda_{i}+\kappa_{i}=0,\\ &a_{i}\geq 0,\forall i,\sum_{i=1}^{d}a_{i}=1,\sum_{i=1}^{d}a_{i}\lambda_{i}\geq m,\sum_{i=1}^{d}a_{i}\lambda_{i}\leq M,\\ &\eta\geq 0,\gamma\geq 0,\kappa_{i}\geq 0,\forall i,\\ &\eta(\sum_{i=1}^{d}a_{i}\lambda_{i}-m)=0,\gamma(M-\sum_{i=1}^{d}a_{i}\lambda_{i})=0,\kappa_{i}a_{i}=0,\forall i.\end{cases}$
Since $\lambda_{i}$’s are distinct, to satisfy the first condition above, $\kappa_{i}$’s can only be zero for at most two indices. Hence, $a_{i}$ can only be non-zero for at most two indices, denoted by $a_{i}$ and $a_{j}$. We write $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$ and hence $\lambda_{i}>\lambda_{j}$. Let $y=\|x\|_{V}^{2}\in[\lambda_{j},\lambda_{i}]$, we have $a\lambda_{i}+(1-a)\lambda_{j}=y$ and $a\lambda_{i}^{-1}+(1-a)\lambda_{j}^{-1}=\|x\|_{V^{-1}}^{2}$.
By direct computation, $\|x\|_{V^{-1}}^{2}=\frac{1}{\lambda_{i}}+\frac{1}{\lambda_{j}}-\frac{y}{\lambda_{i}\lambda_{j}}\eqqcolon f(y,\lambda_{i},\lambda_{j})$, and clearly,
$\begin{cases}\frac{\partial f}{\partial y}&=-\frac{1}{\lambda_{i}\lambda_{j}}<0,\\ \frac{\partial f}{\partial\lambda_{i}}&=(\frac{y}{\lambda_{j}}-1)\frac{1}{\lambda_{i}^{2}}>0,\\ \frac{\partial f}{\partial\lambda_{j}}&=(\frac{y}{\lambda_{i}}-1)\frac{1}{\lambda_{j}^{2}}<0.\end{cases}$
Then the maximum of $\|x\|_{V^{-1}}^{2}$ is obtained by setting $\lambda_{i}=\lambda_{1}$, $\lambda_{j}=\lambda_{d}$, $y=m$, and we have
$x^{*}_{\max}$ $=\frac{m-\lambda_{d}}{\lambda_{1}-\lambda_{d}}\lambda_{1}+\frac{\lambda_{1}-m}{\lambda_{1}-\lambda_{d}}\lambda_{d},\quad\|x^{*}_{\max}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{1}}+\frac{1}{\lambda_{d}}-\frac{m}{\lambda_{1}\lambda_{d}}.$
Similarly, by investigating the KKT conditions for the LP problem that minimize $\sum_{i=1}^{d}a_{i}\lambda_{i}^{-1}$, the minimum of $\|x\|_{V^{-1}}^{2}$ is obtained by some $x=a\lambda_{i}+(1-a)\lambda_{j}$ where $i<j$, $\lambda_{i}>\lambda_{j}$. Setting $\lambda_{u}=\min\{\lambda_{i}\geq M\}$, $\lambda_{l}=\max\{\lambda_{i}\leq M\}$, $y=M$, and we have
$x^{*}_{\min}$ $=\frac{M-\lambda_{l}}{\lambda_{u}-\lambda_{l}}\lambda_{u}+\frac{\lambda_{u}-M}{\lambda_{u}-\lambda_{l}}\lambda_{l},\quad\|x^{*}_{\min}\|_{V^{-1}}^{2}=\frac{1}{\lambda_{u}}+\frac{1}{\lambda_{l}}-\frac{M}{\lambda_{u}\lambda_{l}}.$
This concludes the proof. $\square$
Appendix D Other Proofs
D.1 Proof of Proposition 3
By Propostion 1, we have
${\mathbb{P}}\left[\widehat{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\|\widehat{\theta}_{t}-\theta^{\star}_{t}\|_{V_{t}}\leq\beta^{RLS}_{t}(\delta^{\prime})\right]$
$\geq 1-\frac{\delta}{2}.$
Similarly, by Definition 2, we have
${\mathbb{P}}\left[\widetilde{\mathcal{A}}_{T}\right]$ $={\mathbb{P}}\left[\cap_{t=1}^{T}\left\{\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right\}\right]$
$\geq 1-\sum_{t=1}^{T}{\mathbb{P}}\left[\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})|\mathcal{F}_{t}\right]$
$\geq 1-\frac{\delta}{2}.$
Combining the two inequalities above, we have ${\mathbb{P}}\left[\mathcal{A}_{T}\right]\geq 1-\delta$$\square$
D.2 Proof of Proposition 4
Since ${\theta}^{\star}\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})$ and $\widetilde{\theta}_{t}\in\mathcal{E}_{t}^{PVT}(\delta^{\prime})$, it holds that
$\|{\theta}^{\star}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}^{RLS}(\delta^{\prime}),\quad\|\widetilde{\theta}_{t}-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta^{PVT}_{t}(\delta^{\prime})=\iota_{t}\beta^{RLS}_{t}(\delta^{\prime}).$ (9)
In the following, we drop the dependence on $\delta^{\prime}$ for notation simplicity. By direct computation
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $=\left(\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle\right)+\left(\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\right)$
$\quad+\left(\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle\right)+\left(\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right).$
To bound the second term on the right hand side, recall by Equation (3) the POFUL action $\widetilde{x}_{t}$ satisfies
$\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}\leq\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle+\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}.$
Rearranging the ineuqlity, we obtain $\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle\leq\tau_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}-\tau_{t}\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}$.
The other three terms are bounded similarly using the Cauchy-Schwarz inequality, the triangle ineuqlity of the $V_{t}^{-1}$-norm, and the concentration condition (9). As a result, we have
$\langle{x^{\star}_{t}},{\theta^{\star}}\rangle-\langle{x^{\star}_{t}},{\widetilde{\theta}_{t}}\rangle$ $\leq(1+\iota_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS},$
$\langle{\widetilde{x}_{t}},{\widetilde{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle$ $\leq\iota_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}),$
$\langle{\widetilde{x}_{t}},{\widehat{\theta}_{t}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle$ $\leq\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}(\delta^{\prime}).$
Combing all terms above, we have
$\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\leq(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta^{RLS}_{t}.$
$\square$
D.3 Proof of Theorem 1
We formally prove Theorem 1 for completeness. The proof techniques are developed in previous papers [Abbasi-Yadkori et al., 2011, Agrawal and Goyal, 2013, Abeille et al., 2017].
Throughout the proof, we condition on the event $\mathcal{A}_{T}$, which holds with probability $1-\delta$ by Proposition 3. Applying Proposition 4, we obtain
$\mathcal{R}(T)$ $\leq\sum_{t=1}^{T}\left(\langle{x_{t}^{\star}},{\theta^{\star}}\rangle-\langle{\widetilde{x}_{t}},{\theta^{\star}}\rangle\right)\mathbb{I}\{\mathcal{A}_{t}\}$
$\leq\sum_{t=1}^{T}(1+\iota_{t}-\tau_{t})\|x^{\star}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime})+(1+\iota_{t}+\tau_{t})\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
Recall $\|x^{\star}_{t}\|_{V^{-1}_{t}}=\alpha_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}$ and $\mu_{t}=\alpha_{t}(1+\iota_{t}-\tau_{t})+1+\iota_{t}+\tau_{t}$, we have
$\mathcal{R}(T)\leq\sum_{t=1}^{T}\mu_{t}\|\widetilde{x}_{t}\|_{V^{-1}_{t}}\beta_{t}^{RLS}(\delta^{\prime}).$
Applying the Cauchy-Schwarz inequality and Proposition 2, note that $\max_{t\in[T]}\beta_{t}^{RLS}(\delta^{\prime})=\beta_{T}^{RLS}(\delta^{\prime})$, we obtain
$\mathcal{R}(T)$ $\leq\sqrt{\sum_{t=1}^{T}\|\widetilde{x}_{t}\|_{V_{t}^{-1}}^{2}}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}\left(\beta_{t}^{RLS}(\delta^{\prime})\right)^{2}}\leq\sqrt{2d\log\left(1+\frac{T}{\lambda}\right)}\sqrt{\sum_{t=1}^{T}\mu_{t}^{2}}\beta_{T}^{RLS}(\delta^{\prime}).$
This concludes the proof.$\square$
D.4 Proof of Proposition 5
This corresponds to the case where $m,M\to\lambda_{1}$ in Lemma 3. Without loss of generality, we let $M>\lambda_{2}$. Then $\widehat{\alpha}_{t}=(\lambda_{1}^{-1}+\lambda_{d}^{-1}-m{\lambda_{1}^{-1}\lambda_{d}^{-1}})/(\lambda_{k}^{-1}+\lambda_{k+1}^{-1}-{M}{\lambda_{k}^{-1}\lambda_{k+1}^{-1}})\to\lambda_{1}^{-1}/\lambda_{1}^{-1}=1$$\square$
D.5 Proof of Theorem 3
Define $x^{\dagger}(\theta)\coloneqq\theta/\|\theta\|$. The conclusion of Theorem 2 holds for $x^{\dagger}(\theta)$, i.e. for all $t\in[1,T]$ and all $s\leq t$, we have $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s}$ and $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}$.
For any $\theta$, the feasible optimal action $x^{\star}_{s}(\theta)$ is the point in $\mathcal{X}_{s}$ that has the smallest distance from $x^{\dagger}(\theta)$. Then we have $\|x^{\dagger}(\theta)-x^{\star}_{s}(\theta)\|\leq\rho_{s}$ by the definition of the Euclidean covering radius. Therefore, let $\theta_{0}$ denote the $\theta$ that attains $\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}$, it holds that
$\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}=\|x^{\star}_{s}(\theta_{0})\|_{V_{s}^{-1}}$ $\geq\|x^{\dagger}(\theta_{0})\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\inf_{\theta\in\mathcal{E}_{s}^{PVT}}\|x^{\dagger}(\theta)\|_{V_{s}^{-1}}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$
$\geq\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}).$
Similarly, we have $\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s})$. As a result, it holds that $\alpha_{s}\leq(\Phi_{s}+\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))/(\Psi_{s}-\rho_{s}\lambda_{d}^{-\frac{1}{2}}(V_{s}))$$\square$
Appendix E TS-MR and Greedy-MR Algorithms
Algorithm 2 TS-MR (Greedy-MR)
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
     $\widehat{\mu_{t}}\leftarrow\widehat{\alpha}_{t}(1+\iota_{t})+1+\iota_{t}$
     if $\widehat{\mu}_{t}\leq\mu$ then
        Sample $\eta_{t}\sim\mathcal{D}^{SA}(\delta^{\prime})$ (defined in Section 3)
        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     else
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     end if
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
  end for